STATISTICAL DISTRIBUTION OF DILUTION FACTORS IN SCCl_{2}

\underline{SANDRA LEE} and MARTIN GRUEBELE, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL 61801.

Intramolecular vibrational energy redistribution (IVR) allows for excitation energy to ”flow” to other vibrational modes. The dilution factor is one measure of IVR. Stimulated emission pumping (SEP) spectra at 0.3cm^{−1} resolution from several vibrational states of both the Ã state and B̃ state of thiophosgene (SCCl_{2}) are dumped to the 7,000 cm^{−1} to 10,000 cm^{−1} region on the X̃ state. These SEP spectra are then analyzed using autocorrelation. In addition, simulated spectra are used for comparison of the data against two different models of the statistical distribution of dilution factors.